3^(x^2+x)=9
And if log base 3 (2) =a log base 3 (5) =b write log base 3 (8/25) in terms of a and b

Answer by stanbon(75887)

3^(x^2+x)=9
---
3^(x^2+x) = 3^2
---
x^2+x = 2
x^2+x-2 = 0
(x+2)(x-1) = 0
x = -2 or x = 1
===========
And if log base 3 (2) =a log base 3 (5) =b write log base 3 (8/25) in terms of a and b
---
log3(8/25) = log3(8) - log3(25)
---
= 3log3(2) - 2log3(5)
---
= 3a - 2b
===================
cheers,
Stan H.